OpenStudy (anonymous):

Calculate the following integral \[\int\limits_{0}^{\pi/2} \sin3\theta \cos \theta dx \]

OpenStudy (mimi_x3):

\[ sin 3xcosx = \frac{1}{2} (sin(3x-x) + sin(3x+x) \] \[\frac{1}{2} sin(2x) + sin(4x)\]

OpenStudy (mimi_x3):

I used \(x\) instead of \(\theta\) cos it's easier to type.

OpenStudy (mimi_x3):

\[ \frac{1}{2}\int sin(2x) + \int sin(4x)\]

OpenStudy (mimi_x3):

\[ = \frac{1}{2} * -\frac{1}{2} cos(2x) + -\frac{1}{4}sin(4x) +c\]

OpenStudy (mimi_x3):

don't forget to sub in the limit.

OpenStudy (anonymous):

Thank you, it gives me the right answer